OpenStudy (anonymous):

solve the following system by graphing and identify the point of intersection -5x+2y=-4 2x-y=2

OpenStudy (anonymous):

(0,-2)

OpenStudy (anonymous):

hah!

OpenStudy (anonymous):

-5x+2y=-4 (2x-y=2)x2= 4x-2y=4 -5x+2y=-4 4x-2y=4 x=0,y=-2 Point of intersection in (0,-2)

OpenStudy (anonymous):

Do you need to know how to do it or are you just after an answer?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

im only posting the ones i don't understand

OpenStudy (anonymous):

OK well sattelit posted a correct answer but I will show you step by step how to do it.

OpenStudy (anonymous):

Before I do however have you looked at born2die20 answer?